Find the general sol (including equilibrium and singular sol if there are) of the ODE (all sol need to be real sol.
(c) (15 pts.) y(4)-8y(3) + 16y = 0
Answer #1

\\ \text{The auxiliary equation is }r^4-8r^3+16r^2=0\\ \Rightarrow r^2(r^2-8r+16)=0\\ \Rightarrow r^2(r-4)^2=0\\ \Rightarrow r_1=r_2=0,\quad r_3=r_4=4.\\\\ \text{Then the general solution is }y(t)=c_1+c_2t+(c_3+c_4t)e^{4t}.